OpenStudy (mathlegend):

Solve for h: 2πrh + 2πr^2 = S

OpenStudy (mathlegend):

@Hero

hero (hero):

Did you see my reponse to the previous question you posted?

OpenStudy (mathlegend):

Let me check it.

hero (hero):

Anyways for this problem, subtract \(2πr^2\) from both sides. Let me know what you get afterwards.

OpenStudy (mathlegend):

2πrh=S-2πr^2

OpenStudy (mathlegend):

?

OpenStudy (zzr0ck3r):

right now solve for h

hero (hero):

Okay, good. Now divide both sides by \(2πr\). Let me know what you get

hero (hero):

@zzr0ck3r, he has to understand the process first. saying "now solve for h" probably means nothing to him

OpenStudy (zzr0ck3r):

yeah, im on te phone. I noticed that after I said it....

OpenStudy (mathlegend):

h=S- (I don't know what goes next becuase I'm left with the exponent by itself.)

hero (hero):

Hint: You will be left with a fraction on the right side after dividing both sides by \(2πr\)

hero (hero):

Your left side looks good so far.

OpenStudy (mathlegend):

Yeah but I get... \[\frac{ 2πr^2 }{ 2πr }\]

hero (hero):

What happened to the S ?

hero (hero):

That doesn't disappear

OpenStudy (mathlegend):

S-\[\frac{ 2πr^2 }{ 2πr}-S\]

hero (hero):

You have to divide the entire right side by \(2πr\), not just \(2πr^2\)

OpenStudy (mathlegend):

ok

OpenStudy (mathlegend):

\[\frac{ S-2πr^2 }{ 2πr}\]

hero (hero):

\[h = \frac{S - 2\pi r^2}{2 \pi r}\]

hero (hero):

Good job

OpenStudy (mathlegend):

Thanks @Hero

hero (hero):

Do you want to discuss the other question?

OpenStudy (mathlegend):

Yeah.

OpenStudy (mathlegend):

A=2πpw

hero (hero):

I pinged you to the question